### v1

# Quantum noise spectroscopy of superconducting critical dynamics and vortex fluctuations in a high-temperature cuprate

Abstract
Characterizing the low-energy dynamics of quantum materials is crucial to our understanding of strongly correlated electronic states. However, extracting universal dynamical features requires resolving correlations at both low energy and momentum. Here, we introduce nitrogen-vacancy (NV) centers in diamond as a novel and powerful quantum sensing platform of superconducting materials. We demonstrate the strengths of our approach by probing several low-energy phenomena in high-$T_{c}$ cuprate Bi2Sr2CaCu2O8+δ (BSCCO) — gapless quasiparticle excitations, critical fluctuations at the metal-superconductor transition and kinetics of vortices. In the absence of an applied magnetic field, we find a sharp reduction in the NV relaxation time ($T_{1}$) near the critical temperature $T_{c}\approx 90~{}$K, attributed to supercurrent-fluctuation induced magnetic noise. Crucially, the temperature-scaling of the noise near criticality deviates from the Bardeen–Cooper–Schrieffer (BCS) mean-field prediction and reflects critical order parameter fluctuations, allowing us to determine both static and dynamical critical exponents for the transition. When a small field is applied, we detect a broad and asymmetric reduction of $T_{1}$ near $T_{c}$, indicating significant field-induced smearing of the transition. By analyzing the scaling of the BSCCO-induced relaxation rate with the field strength, we unveil evidence in favor of a vortex liquid phase. Finally, deep inside the superconducting phase, we employ NV decoherence ($T_{2}$) spectroscopy to observe strong magnetic fluctuations in the low-frequency regime, suggesting the presence of complex vortex-solid fluctuations. Our results establish NV-based noise spectroscopy as a versatile platform for probing dynamical phenomena in superconductors, with frequency and length scales complementary to existing techniques.
Introduction — Strong electronic interactions are responsible for a plethora of complex emergent phenomena observed in quantum materials. This is perhaps best exemplified by unconventional superconductors, where strong electronic interactions induce a macroscopic quantum coherence of electron-pairs and dissipationless flow of currents tinkham ; keimer2015quantum ; zhou2021high . Key features of this state are encoded in its low-energy dynamics, manifesting as quasiparticle excitations arising from the breaking of electron pairs bardeen1957microscopic ; bardeen1957theory ; bogoljubov1958new , slow collective modes such as critical pairing fluctuations near the metal–superconductor phase transition larkin2005theory ; schmid1966time ; maki1968critical ; cyrot1973ginzburg ; schuller2006time , or topological excitations such as vortices carrying quantized magnetic flux tinkham ; blatter1994vortices ; halperin1979resistive ; Beasley ; Nelson_VL . Over the years, several experimental techniques have been developed to probe these universal, low-energy features comin2016resonant ; kaindl2000ultrafast ; de2021colloquium ; parker:2010 ; essaleh:2018 ; hore:2015 .
Figure 1: NV noise spectroscopy of superconducting dynamics. (a) Summary of different experimental techniques for probing superconductivity, organized by their frequency and lengthscales, highlighting the complementarity of our qubit-based noise spectroscopy (NV $T_{1}$ and $T_{2}$) and DC magnetometry barry:2020b . Techniques depicted include: electrical transport essaleh:2018 , superconducting quantum interference device (SQUID) kirtley:1999 ; dusad:2019 , Lorentz microscopy harada:1992 ; silhanek:2021 , X-ray/neutron scattering comin2016resonant ; muhlbauer:2019 , magnetic resonance with nuclear, electronic, or muonic spins (NMR/ESR/µSR) hore:2015 ; hillier:2022 , angle-resolved photoemission spectroscopy (APRES) liu2008development ; kitamura2022development , low-energy and transmission electron microscopy (LEEM/TEM) altman:2010 ; moradifar:2023 , and scanning tunneling microscope (STM) parker:2010 . Some techniques, while capable of accessing high momentum data, require macroscopic sample volumes and, thus, are only suitable for probing mm-scale samples. Techniques like pump-probe spectroscopy kaindl2000ultrafast ; de2021colloquium , which are known to drive the sample strongly out of equilibrium, are not included. (b) Schematic of the experimental set-up: a thin film ($\sim 200$ nm thickness) of BSCCO is directly exfoliated onto a diamond plate containing an ensemble of NV centers within a thin layer ($\sim 50$ nm) beneath its surface. Current fluctuations in the BSCCO sample induce magnetic field noise at the NV locations which can be detected by either depolarization ($T_{1}$) or decoherence ($T_{2}$) noise spectroscopy. An out of plane magnetic field $H$ is applied, which is parallel to the quantization axis of the [111] NV group. The NV exhibits a splitting of its $|m_{s}=\pm 1\rangle$ sublevels, which is determined by the local magnetic field $B_{z}$, centered around a zero-field splitting $D_{\mathrm{gs}}=2.87~{}$GHz. (c) Schematic H-T phase diagram of a typical type-II superconductor tinkham ; keimer2015quantum ; zhou2021high . The green arrows indicate zero-field cooling ($H=0$) and field-cooling ($H>0$) pathways used in our experiments.
Broadly, techniques for measuring correlated quantum dynamics may be categorized into scattering methods, local spectroscopic probes, and bulk measurements [Fig. 1(a)]. Scattering probes (e.g. X-ray scattering comin2016resonant ) provide both energy and momentum-resolved data, but they typically require experimental apparatus beyond table-top settings, large sample volumes, and can drive the system out of equilibrium kaindl2000ultrafast ; de2021colloquium . Local probes such as scanning tunneling microscopy (STM) parker:2010 are applicable to low-dimensional samples and provide lattice-scale spatial resolution, yet they exhibit limited momentum resolution and typically operate at low temperatures. By contrast, bulk measurements (including transport essaleh:2018 and nuclear magnetic resonance hore:2015 ) can operate over a wide temperature range, but generally demand large sample volume. Therefore, resolving spectral features at long wave-lengths below the meV scale in two-dimensional materials and thin film samples remains a challenge, leading to an unfulfilled gap for experimental techniques that can probe equilibrium dynamics at both low-energies and a variety of length-scales.
In this work, we fill this gap by introducing a non-invasive, table-top experimental technique — quantum noise spectroscopy — to probe superconducting dynamics with simultaneous spatial and temporal resolution. Specifically, by using an ensemble of nitrogen-vacancy (NV) centers positioned in close proximity to a thin film high-$T_{c}$ cuprate — Bi2Sr2CaCu2O8+δ (BSCCO, Bi-2212), we perform noise spectroscopy of superconducting fluctuations as a function of temperature, frequency and applied magnetic field. We demonstrate in-situ detection of spatially resolved dynamics spanning a broad frequency range (MHz to GHz) and length-scales (nm to $\mathrm{\mu m}$), thus offering a powerful complementary toolset to existing techniques for probing superconductivity [Fig. 1(a)].
Our main results are threefold. First, at zero applied field, we study the fluctuations arising from both long-lived quasiparticle excitations at low temperature, and collective critical modes near the metal-superconductor transition temperature $T_{c}$. These fluctuations induce GHz-scale magnetic noise at the location of the nearby NVs which can be characterized via NV depolarization ($T_{1}$) spectroscopy. Leveraging this capability, we quantitatively measure the characteristic time-scale $\tau_{0}$ of critical slowing down of pairing fluctuations near $T_{c}$, as well as the static and dynamical critical exponents associated with the phase transition landau2013statistical ; landau1980statistical ; hohenberg1977theory .
Second, we directly probe the broadening of the transition in the presence of a small applied field $H$. At non-zero $H$, the system hosts superconducting vortices whose dynamics induce additional magnetic field fluctuations. The field dependence of the NV relaxation rate is consistent with the existence of an intermediate-temperature vortex liquid phase blatter1994vortices ; halperin1979resistive , where magnetic flux lines diffuse through the superconducting film.
Finally, we employ decoherence ($T_{2}$) noise spectroscopy to study lower-frequency fluctuations in the MHz regime machado2023quantum . Deep within the superconducting phase, we observe a distinct noise mechanism that significantly enhances the NV decoherence rate. These strong, low-frequency magnetic fluctuations potentially arise from thermally assisted flux motion in a vortex solid phase blatter1994vortices ; MFisher89 ; fisherfisherhuse .
Figure 2: Observing Meissner effect and superconducting vortices. (a) ODMR spectrum of NV centers under BSCCO with an applied field $H=15$ G at temperatures from 10 K to 120 K. The colors represent the normalized fluorescence contrast (FL) of the spectrum. The gray dashed line marks the superconducting transition temperature $T_{c}\approx 90$ K. Top: at 10 K, ODMR spectrum of NV centers under BSCCO at $H=0$ (yellow), where the intrinsic splitting $2\Pi_{\perp}$ originates from the local electric field mittiga2018imaging . At $H=15$ G, the spectrum (blue) remains largely unchanged, indicating a strong magnetic field repulsion due to the Meissner effect. Bottom: when the bias field is not suppressed (120 K), the [111] NV group aligned with the bias field exhibits a splitting $\Delta$, while the remaining NVs in other three crystal orientations are degenerate, contributing to the inner two peaks. (b) Optical image of the thin film BSCCO exfoliated on a diamond surface. (c) Local magnetic field $B_{z}$ as a function of the applied field $H$ along a line-cut across the BSCCO [black dashed line in (b)] at 30 K. Compared to NV centers far away from BSCCO, those under BSCCO (gray region) experience smaller local magnetic fields due to the Meissner suppression. Near the edge of the BSCCO, $B_{z}$ gets enhanced due to the concentration of the magnetic field lines. (d) Schematic of magnetic flux penetrating a superconductor in its vortex phase. Top: local field image of a defect-trapped superconducting vortex at $H=40$ G and $T=40$ K. (e) Schematic of magnetic flux being repelled by a superconductor in its Meissner phase.
Experimental setup— For our experiments, we choose Bi2Sr2CaCu2O8+δ (Bi-2212), a prominent type-II superconductor with a layered crystal structure and a high critical temperature $T_{c}$ exceeding the boiling point of liquid nitrogen keimer2015quantum ; zhou2021high ; blatter1994vortices . This compound exhibits a rich and complex H-T phase diagram, making it an ideal platform for investigating a variety of different types of low-energy dynamics [Fig. 1(c)]. Bulk crystals of nearly optimally doped BSCCO-2212 are grown by the traveling floating zone method wen2008large , with $T_{c}\approx 90~{}$K measured via transport (see Methods). To perform noise spectroscopy with NV centers, we exfoliate the BSCCO to a 200200200~{}200nm thick film and transfer it directly onto a diamond sample containing an ensemble of NV centers in a thin layer $\sim 50$ nm beneath the diamond surface [Fig. 1(b)].
Figure 3: Zero-field superconducting fluctuations. (a) $T_{1}$ decay profile for NV centers located under BSCCO (left, blue) and away from BSCCO (right, orange) at 110, 93 and 40 K. Inset: Location of the measured NV centers with respect to the BSCCO sample. (b) The $T_{1}$ relaxation timescales as a function of temperature measured on the NV centers located under and away from BSCCO. (c) Spin relaxation rate, $1/T_{1}$, along a spatial line-cut across the BSCCO sample [see (a) inset] at 110, 93 and 40 K. The relaxation rate increases for NV centers under BSCCO (shadow region), particularly near the critical temperature $T_{c}$. (d) The BSCCO-induced decay rate of NVs, $\Gamma_{1}(T)$, as a function of temperature with $H=0$. Inset: a power-law relationship at low temperatures: $\Gamma_{1}(T)\propto T^{2}$, is consistent with a mean-field approach accounting for nodal quasiparticles in d-wave superconductors. (e) $\Gamma_{1}(T)$ as a function of the absolute temperature deviation from the critical temperature, $|T-T_{c}|$, for both the metallic and superconducting (SC) sides of the transition. BCS mean-field quasiparticle dynamics (dashed line) predicts $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$. However, our experimental data is consistent with $x\approx 1$, suggesting that additional superconducting order parameter fluctuations are important. This data aligns closely with a theoretical model of critical Langevin dynamics of the order parameter (solid lines).
Each NV center hosts a spin-1 electronic ground state, with Hamiltonian rovny2024nanoscale :
$\begin{split}\mathcal{H}=&~{}D_{\mathrm{gs}}S_{z}^{2}+\gamma_{e}B_{z}S_{z}\\ &+\Pi_{x}(S_{y}^{2}-S_{x}^{2})+\Pi_{y}(S_{x}S_{y}+S_{y}S_{x}),\end{split}$ (1)
where {$S_{x}$, $S_{y}$, $S_{z}$} are the spin-1 operators with the N–V axis defining the spin-quantization axis ($\hat{z}$), $D_{\mathrm{gs}}=2.87~{}$GHz is the zero-field splitting between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sublevels, $\Pi_{\perp}=\sqrt{\Pi_{x}^{2}+\Pi_{y}^{2}}=6.6~{}$MHz characterizes the coupling to local electric fields mittiga2018imaging , $B_{z}$ is the local magnetic field along the NV axis, and $\gamma_{e}$ is the electronic spin gyromagnetic ratio.
The NV energy levels can be probed via optically detected magnetic resonance (ODMR) spectroscopy where one detects a decrease in NV fluorescence intensity when sweeping the frequency of an applied microwave through the resonances $|m_{s}=0\rangle\leftrightarrow|m_{s}=\pm 1\rangle$. In the experiment, we use a [111]-cut diamond crystal and apply a tunable magnetic field $H$ applied along the out-of-plane direction [Fig. 1(b)]. In this geometry, the NV crystallographic group aligned with the applied field exhibits the largest splitting, $\Delta=\sqrt{(2\Pi_{\perp})^{2}+(2\gamma_{e}B_{z})^{2}}$ between $|m_{s}=\pm 1\rangle$ sublevels, where $B_{z}$ includes both contributions from the applied field $H$ and the $B^{\mathrm{s}}_{z}$ generated from the superconducting sample ($B_{z}=H+B_{z}^{\mathrm{s}}$). The other three NV groups are always observed to be degenerate.
Imaging the Meissner effect and flux vortices — Using the NV centers as sensitive local magnetometers, we begin by determining the presence of superconductivity in BSCCO via the DC Meissner effect schlussel2018wide ; yip2019measuring ; lesik2019magnetic ; bhattacharyya2023imaging ; nusran2018spatially ; ho2024studying . After zero-field cooling the sample down to 101010~{}10K (well below $T_{c}$), we perform an ODMR measurement with $H=0$ [Fig. 2(a)]. The resulting spectrum shows two resonances separated by the intrinsic splitting $2\Pi_{\perp}$. When we then apply a small field $H=15~{}$G, the ODMR spectrum remains largely unchanged, displays only a slightly larger splitting, suggesting that the magnetic field generated by the sample, $B_{z}^{\mathrm{s}}\approx-H$. Indeed, the extracted local field at the NV, $B_{z}=2.5~{}$G $\ll H=15~{}$G, corroborating the presence of superconducting Meissner phase that expels the external field. We emphasize that the observation of a near-complete cancellation of $H$ underscores the close proximity of our NV sensors to the superconducting BSCCO.
As we increase the temperature, the near-complete expulsion of the external field persists $T\sim 40~{}$K. Beyond this point, ODMR spectrum splits into a characteristic four resonances, indicating that the field is no longer fully canceled and that the system has transitioned into a vortex phase, allowing magnetic flux to penetrate the sample. As the temperature further rises, the vortex density increases, raising the local field, $B_{z}$ until it reaches the applied field $H$ at $T=T_{c}\approx 90~{}$K, Interestingly, right below $T_{c}$, the local field $B_{z}$ is measured to be slightly greater than $H$, which may be attributed to a paramagnetic Meissner effect geim1998paramagnetic .
Leveraging the local nature of our sensor, we then investigate the spatial dependence of the Meissner suppression. Performing ODMR spectroscopy along a line-cut across the BSCCO [black dashed line in Fig. 2(b)], we directly extract the local magnetic field $B_{z}$ as a function of spatial position at $T=30$ K [Fig. 2(c)]. Beneath the sample, the Meissner suppression is clearly observed at small applied field $H\lesssim 20~{}$G. However, when $H$ exceeds 25 G, which is on the same order as the independently characterized lower critical field $H_{\mathrm{c1}}$ (see Methods), the Meissner effect is suppressed, and $H$ is only partially expelled. Crucially, the residual field also exhibits spatial variations across the sample, indicating that magnetic flux penetrates the BSCCO, forming superconducting vortices. One such vortex, possibly pinned by a local defect, is imaged in Figure 2(d), where the locally measured magnetic field $B_{z}$ exceeds the applied field $H=40~{}$G due to the bunching of magnetic field lines.
Zero-field superconducting fluctuations — Employing the full capabilities of the NV center, we go beyond conventional static magnetic field measurements by performing $T_{1}$ relaxometry of the NV to study superconducting fluctuations [Fig. 1(a)]. To this end, we first optically polarize the NV spin state to $|m_{s}=0\rangle$ and subsequently measure the timescale $T_{1}$ over which the spin relaxes back to thermal equilibrium [Fig. 3(a)]. Although the intrinsic relaxation of NV centers is primarily driven by spin-phonon interactions within the diamond itself, additional magnetic fluctuations from BSCCO, resonant with the NV’s bare frequency ($D_{\mathrm{gs}}=2.87~{}$GHz), can expedite the spin relaxation process and reduce the $T_{1}$ of NV.
This is indeed borne out by our data. As shown in Figures 3(a)(b), NV centers located beneath the BSCCO exhibit shorter $T_{1}$ compared to NV centers situated away from the BSCCO across the entire temperature range covering the superconducting phase, criticality and metallic phase. This trend is further highlighted in Figure 3(c), which shows that the NV relaxation rate ($1/T_{1}$) as a function of a spatial line-cut across three distinct temperature regimes, with faster relaxation rates observed in all cases for NV centers beneath the BSCCO sample.
To quantitatively analyze the observed magnetic noise, we extract the BSCCO-induced decay rate, $\Gamma_{1}(T)$, by subtracting the intrinsic relaxation rate — measured from NV centers located far from the BSCCO sample — from the total relaxation rate of NV centers situated beneath the BSCCO [Fig. 3(d)]. At high temperatures, $T>T_{c}$, $\Gamma_{1}(T)$ is finite due to Johnson-Nyquist noise in the metallic phase kolkowitz2015probing ; agarwal2017magnetic ; hsieh2019imaging . At very low temperatures, $T\ll T_{c}$, the magnetic noise is significantly suppressed in the superconductor. Crucially, the form of this suppression offers insights into the nature of the pairing symmetry. While in conventional s-wave superconductors $\Gamma_{1}(T)$ is exponentially suppressed at low temperatures due to a fully gapped Bogoliubov quasiparticle spectrum  bardeen1957theory ; bardeen1957microscopic ; chatterjee2022single ; dolgirev2022characterizing ; kelly2024superconductivity ; li2024observation , the presence of nodal quasiparticles in d-wave layered superconductors implies a distinct power-law suppression of $\Gamma_{1}(T)\propto T^{2}$ dolgirev2022characterizing ; chatterjee2022single . Indeed, this expectation is consistent with our observations [Fig. 3(d), inset], corroborating the nodal d-wave nature of superconductivity in BSCCO keimer2015quantum .
The most striking feature of Fig. 3(d) is the sharp and symmetric peak in $\Gamma_{1}$ near the critical temperatures $T_{c}$. This peak arises from strong fluctuations in the superconducting order parameter, resulting in enhanced current fluctuations that act as source of magnetic noise at the metal-superconductor phase transition. Crucially, this enhancement is observed across the entire BSCCO sample, as emphasized by the spatial line cut near $T_{c}$, Fig. 3(c). Interestingly, near $T_{c}$, $\Gamma_{1}(T)$ begins to show marked deviations from mean-field expectations for d-wave superconductors. In particular, mean-field theory predicts a divergence of $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$ up to logarithmic corrections dolgirev2022characterizing . In contrast, the experimentally measured $\Gamma_{1}$ at criticality is distinctly more singular, with $x\approx 1$ [Fig. 3(e)]. This deviation arises because BCS theory neglects both amplitude and phase fluctuations of the superconducting order parameter $\psi$, which play a significant role near the critical point emery1995importance ; maki1968critical . These effects are particularly important for thin-film superconductors where fluctuations are expected to be stronger tinkham . Consequently, quantitative characterizing the divergent time-scale associated with dissipative dynamics of the superconducting order parameter near $T_{c}$ remains an important open question, which we turn to next.
To properly account for low-energy critical fluctuations, we consider the following phenomenological Langevin dynamics of the superconducting order parameter $\psi(\mathbf{r},t)$.
$\partial_{t}\psi(\mathbf{r},t)$ $=-\gamma\frac{\delta F}{\delta\Psi^{*}(\mathbf{r},t)}+\eta(\mathbf{r},t),$
[…]
where $F[\psi]$ is the Ginzburg-Landau free energy functional landau1980statistical ; cyrot1973ginzburg ; schuller2006time ; tinkham , $\eta$ is local white noise that captures the effect of coarse-graining, and $\gamma$ sets the rate of relaxation towards equilibrium hohenberg1977theory . Crucially, by setting $r(T)=a(T-T_{c})$ with $a>0$, we can analyze fluctuations of the order parameter about the minima of the free energy: $\langle\psi\rangle=0$ for $T>T_{c}$ (metallic) and $\langle\psi\rangle=\psi_{0}$ for $T<T_{c}$ (superconducting). Such fluctuations in turn leads to enhanced current-induced magnetic noise at the NV frequency $D_{\mathrm{gs}}$, enhancing the relaxation rate $\Gamma_{1}(T)$ near $T_{c}$. This enhancement of the noise spectral weight at low frequency is a consequence of critical slowing down of order parameter fluctuations: for weak-coupling superconductors it is quantified by the diverging timescale $\tau_{\rm GL}=8\hbar/(k_{B}|T-T_{c}|)$ larkin2005theory ; schmid1966time near $T_{c}$.
We find that our theoretical model (see Methods) can faithfully reproduce both the magnitude of the magnetic noise, and the divergence of $\Gamma_{1}(T)$ near $T_{c}$ with the experimentally observed power-law $x=1$ on both sides of the transition [Fig. 3(e)]. In addition, by fitting $\Gamma_{1}(T)$ with our model, we extract a quantitative time-scale for critical slowing down $\tau_{0}^{\rm fitted,M}\approx 1.2\,\tau_{\rm GL}$ on the metallic side, and $\tau_{0}^{\rm fitted,SC}\approx 0.2\,\tau_{\rm GL}$ on the superconducting side. Remarkably, our results indicate that the decay rates of the order parameter near criticality are comparable to the analytical weak-coupling results $\tau_{0}^{\rm wc,M}=\tau_{\rm GL}$ larkin2005theory on the metallic side and $\tau_{0}^{\rm wc,SC}=0.5\,\tau_{\rm GL}$ on the superconducting side schmid1966time ; KimDiessel , even though high $T_{c}$ cuprates like BSCCO are expected to be strong-coupling superconductors lee2006doping .
Further, our modeling of critical noise also allows us to extract the static critical exponent $\nu$ that determines the divergence of the order parameter correlation length $\xi(T)\propto|T-T_{c}|^{-\nu}$, and the dynamical critical exponent $z$ that relates the correlation time $\tau_{0}$ and correlation length as $\tau_{0}(T)\propto\xi^{z}$. The close agreement of the experimental observations with Langevin dynamics of $\psi$ implies that $\nu=1/2$, and $\tau_{0}(T)\propto|T-T_{c}|^{-1}\propto\xi^{2}$ implies $z=2$, consistent with expectations for a non-conserved order parameter hohenberg1977theory .
Figure 4: In-field criticality and vortex dynamics. (a) BSCCO-induced relaxation rate, $\Gamma_{1}(T)$ (normalized), as a function of temperature at $H=0,40,200$ G. Inset: the measured peak magnitude of $\Gamma_{1}$ as a function of applied field, which is consistent with a linear dependence on the field strength $H$ (dashed line). (b) BSCCO-induced decoherence rate, $\Gamma_{2}(T)$, as a function of temperature at $H=55,120,200$ G. Inset: The measured $T^{\mathrm{XY8}}_{2}$ decoherence timescale as a function of temperature for NVs under and away from BSCCO at $H=200$ G. (c) Schematic of the noise spectrum arising from different dynamical processes in BSCCO. The shaded area marks the most sensitive frequencies of NV relaxation ($\Gamma_{1}$) spectroscopy and decoherence ($\Gamma_{2}$) spectroscopy. Since the NV’s intrinsic decoherence rate far exceeds its relaxation rate, $\Gamma_{2}$ is less sensitive to weak magnetic noise than $\Gamma_{1}$. (d) Left: a trapped vortex imaged in BSCCO under $H=40$ G at 50 K. Right: the measured relaxation time $T_{1}$ and decoherence time $T^{\mathrm{XY8}}_{2}$ of NV centers at different spatial positions across the vortex.
In-field criticality and vortex dynamics — Next we explore how an out-of-plane magnetic field impacts critical fluctuations. Specifically, we apply a small bias magnetic field and perform $T_{1}$ noise spectroscopy at two distinct fields, $H=40$ G and $H=200$ G. Intriguingly, compared with the zero-field $\Gamma_{1}$, two distinct new features emerge [Fig. 4(a)]: (i) The peak in the measured noise, $\Gamma_{1}$, is shifted towards a slightly lower temperature. (ii) While the noise signal approaching criticality from the metallic side remains relatively sharp, from the superconducting side, it increases much more gradually and smoothly towards criticality, resulting in a broad and asymmetric spectrum. The former is expected, as a perpendicular field hinders phase coherence and lowers $T_{c}$. The latter can be attributed to a significantly larger fluctuation regime, dominated by the two-dimensional XY critical behavior Beasley ; halperin1979resistive ; curtis2024probing . Below $T_{c}$, an intermediate vortex liquid phase blatter1994vortices ; fisherfisherhuse ; Nelson_VL is expected to form. In this phase, the motion of the superconducting vortices, which trap magnetic flux quanta, can generate strongly fluctuating magnetic fields at the NV, resulting in enhanced noise.
Further evidence for this scenario can be gleaned from the scaling of the peak magnitude of $\Gamma_{1}$ with the magnetic field $H$ [inset, Fig. 4(a)], where the density of vortices $n_{v}$ increases with increasing field strength. Within a simple model of diffusive vortex motion (see Methods), the local magnetic noise, measured via $\Gamma_{1}(H)$, scales as $D_{v}^{-1}$ where $D_{v}\sim v^{2}\tau_{v}$ is the vortex diffusion constant. We expect the vortex scattering time $\tau_{v}$ to scale inversely proportional to the vortex density $n_{v}$, and the mean-square speed $\langle v^{2}\rangle$ is primarily determined by thermal energy. Consequently, the measured magnetic noise scales as $\Gamma_{1}(H)\propto n_{v}\propto H$, in line with our observation [inset, Fig. 4(a)]. As the temperature decreases further, the vortices either freeze into a lattice or get pinned by defects, resulting in a decrease of $\Gamma_{1}(T)$. This conclusion is consistent with our measurement of noise over a line-cut on a large pinned vortex at low temperatures shown in Fig. 4(d) — $T_{1}$ does not change appreciably across the vortex, indicating that pinned vortices do not contribute to magnetic noise in the GHz frequency range.
$T_{2}$ noise spectroscopy of slow vortex dynamics — At low temperatures, one enters deep into the superconducting phase and the vortex motion freezes out. Nevertheless, with $H>H_{c1}$, BSCCO can exhibit complex vortex dynamics, such as thermally assisted flux jumps between different pinning centers, or de-pinning of vortices to exit the superconducting sample, potentially leading to a revival of magnetic noise. However, our data for $\Gamma_{1}(T)$ in the low temperature regime exhibits no additional features at finite $H$. One possibility is that, unlike the metal-superconducting transition, fluctuations arising from vortex dynamics at low temperatures occur at a much lower frequency range, and therefore eludes detection by $T_{1}$ spectroscopy that is only sensitive to the noise spectrum in the GHz range.
To this end, we perform NV decoherence ($T_{2}$) noise spectroscopy to capture the magnetic noise spectrum in the MHz frequency range machado2023quantum ; xue2024signatures ; ziffer2024quantum ; li2024critical . Specifically, after preparing the NV center into a quantum superposition state, $(|m_{s}=0\rangle+|m_{s}=-1\rangle)/\sqrt{2}$, we apply an advanced dynamical decoupling sequence (XY-8, see Supplementary Information) to measure the decoherence rate induced by BSCCO, $\Gamma_{2}(T)$, as a function of both temperature and bias external field $H$ [Fig. 4(b)]. Remarkably, $\Gamma_{2}(T)$ exhibits a broad noise spectrum with several prominent features. (i) The onset temperature for $\Gamma_{2}(T)$ progressively decreases with increasing field $H$, and does not coincide with the peak of $\Gamma_{1}$ near criticality. (ii) For a fixed $H$, $\Gamma_{2}(T)$ reaches its maximum value at a much lower temperature $T^{*}$ than $\Gamma_{1}(T)$; however, $H$ is larger than the independently measured $H_{c1}$ of the sample at $T^{*}$ (see Methods), indicating that vortices are still present in the sample. (iii) The maximum value of $\Gamma_{2}(T)$ is nearly independent of the external field $H$.
To understand this behavior, we begin by noting that the decoherence time $T_{2}$ on and off BSCCO is nearly identical in both the metallic phase and near criticality [inset of Fig. 4(b)]. This observation indicates that $T_{2}$ noise spectroscopy is less sensitive than $T_{1}$ in detecting both Johnson-Nyquist noise and critical current fluctuations. This difference in sensitivity can be attributed to the intrinsic timescale gap between $T_{1}$ and $T_{2}$. Specifically, the NV $T_{1}$ is around three orders of magnitude larger than $T_{2}$, offering a much longer time window to accumulate the noise signals (see Methods). Consequently, a significantly stronger noise amplitude from BSCCO is required to induce a measurable $T_{2}$ in NV sensors.
Furthermore, $T_{2}$ measured across a line-cut intersecting a pinned vortex show no position dependence, indicating that frozen vortices also do not contribute to magnetic noise in the MHz frequency range. Combining these observations, we deduce that $T_{2}$ is most sensitive to vortex motion within the vortex solid phase MFisher89 ; fisherfisherhuse . The substantial noise magnitude suggests that it likely originates from thermal fluctuations, which can cause positional jumps of fluxes or flux bundles between different pinning centers, resulting in sudden large changes in the local magnetic field. These dynamics are relatively slow, extending only up to the MHz range. Thus, they are not reflected in $\Gamma_{1}$, but can be directly captured by $\Gamma_{2}$ [Fig. 4(c)]. Only at very low temperatures, approximately corresponding to the Meissner phase, does such flux motion begin to freeze out, leading to a decrease in $\Gamma_{2}(T)$ below $40$ K.
Summary and Outlook — For the first time, we demonstrated the use of NV centers as nano-scale quantum sensors to probe three distinct aspects of low-energy dynamics in high-$T_{c}$ superconductors: gapless quasiparticle excitations at low temperatures, critical superconducting fluctuations near $T_{c}$, and vortex motion in an intermediate temperature range in presence of an external magnetic field. The nanoscale proximity and high sensitivity of our NV sensors, combined with their operational frequency range spanning three decades (from MHz to GHz) and sub-micron spatial resolution, enable the measurement of local dynamics beyond the capabilities of conventional techniques [Fig. 1(a)].
Looking forward, our work opens the door to several intriguing directions. First, while our current studies involve NV center ensembles with a spread of sensor-sample distances, an immediate future direction is to precisely adjust the distance $d$ between the quantum sensor and the target material. This can be achieved by using NVs on a scanning tip maletinsky2012robust ; degen2008scanning ; pelliccione2016scanned , or by creating spin sensors in a given layer of Van der Waals materials such as hexagonal boron nitride gottscholl2020initialization ; gong2023coherent ; stern2022room ; gong2023isotope . By tuning the distance $d$, we can adjust the momentum filter function to enable controlled probing over a variety of length scales (from nanometers to micrometers)  machado2023quantum , which is crucial for imaging inhomogeneous dynamics in superconductors bhattacharyya2023imaging ; Dailledouze2025 .
Moreover, our measurement of the critical slowing down time-scale $\tau_{0}$ places strong constraints on theoretical models of the transition and provides new opportunities to study the mechanisms for order-parameter relaxation near criticality. In fact, a reasonable agreement with the weak coupling result $\tau_{0}=\tau_{\rm GL}$ on the metallic side larkin2005theory , but a stronger deviation on the superconducting side schmid1966time , indicates the necessity for further microscopic investigations of pair dynamics in unconventional superconductors.
Finally, our technique is not limited to BSCCO, and can be readily extended to a variety of superconducting materials. For instance, certain superconductors such as hydrides pickard2020superconducting ; drozdov2015conventional and nickelates sun2023signatures ; wen2024probing undergo superconducting transitions under high pressure. Thus, moving beyond temperature-driven transitions, the integration of NV centers into high-pressure diamond anvil cells offers a pathway to study how pressure influences critical fluctuations and vortex dynamics hsieh2019imaging ; lesik2019magnetic ; yip2019measuring ; bhattacharyya2023imaging ; Dailledouze2025 .
Acknowledgements: We gratefully acknowledge discussions with Pavel Dolgirev, Chris Laumann, Peng Cai, Panyu Hou, Srinivas Mandyam, and Zhipan Wang. This work is supported by NSF NRT LinQ 2152221, NSF ExpandQISE 2328837, and the Center for Quantum Leaps at Washington University. O. K. D. and F. M. acknowledge support from the NSF through a grant for ITAMP at Harvard University. The work at Brookhaven National Laboratory is supported by the US Department of Energy, office of Basic Energy Sciences, contract No. DOE-sc0012704. E. A. H. and X. D. acknowledge support by the Gordon and Betty Moore Foundation, grant DOI 10.37807/gbmf11560. S. R., Q. X., and Z. R. acknowledge support under National Science Foundation (NSF) Division of Materials Research Award DMR-2236528.

### v2

@@
-Quantum noise spectroscopy of superconducting critical dynamics and vortex fluctuations in a high-temperature cuprate
+Quantum noise spectroscopy of superconducting dynamics in thin film Bi2Sr2CaCu2O8+δ
 Abstract
 Characterizing the low-energy dynamics of quantum materials is crucial to our understanding of strongly correlated electronic states.
-However, extracting universal dynamical features requires resolving correlations at both low energy and momentum.
-Here, we introduce nitrogen-vacancy (NV) centers in diamond as a novel and powerful quantum sensing platform of superconducting materials.
-We demonstrate the strengths of our approach by probing several low-energy phenomena in high-$T_{c}$ cuprate Bi2Sr2CaCu2O8+δ (BSCCO) — gapless quasiparticle excitations, critical fluctuations at the metal-superconductor transition and kinetics of vortices.
-In the absence of an applied magnetic field, we find a sharp reduction in the NV relaxation time ($T_{1}$) near the critical temperature $T_{c}\approx 90~{}$K, attributed to supercurrent-fluctuation induced magnetic noise.
-Crucially, the temperature-scaling of the noise near criticality deviates from the Bardeen–Cooper–Schrieffer (BCS) mean-field prediction and reflects critical order parameter fluctuations, allowing us to determine both static and dynamical critical exponents for the transition.
-When a small field is applied, we detect a broad and asymmetric reduction of $T_{1}$ near $T_{c}$, indicating significant field-induced smearing of the transition.
-By analyzing the scaling of the BSCCO-induced relaxation rate with the field strength, we unveil evidence in favor of a vortex liquid phase.
-Finally, deep inside the superconducting phase, we employ NV decoherence ($T_{2}$) spectroscopy to observe strong magnetic fluctuations in the low-frequency regime, suggesting the presence of complex vortex-solid fluctuations.
-Our results establish NV-based noise spectroscopy as a versatile platform for probing dynamical phenomena in superconductors, with frequency and length scales complementary to existing techniques.
-Introduction — Strong electronic interactions are responsible for a plethora of complex emergent phenomena observed in quantum materials.
-This is perhaps best exemplified by unconventional superconductors, where strong electronic interactions induce a macroscopic quantum coherence of electron-pairs and dissipationless flow of currents tinkham ; keimer2015quantum ; zhou2021high .
-Key features of this state are encoded in its low-energy dynamics, manifesting as quasiparticle excitations arising from the breaking of electron pairs bardeen1957microscopic ; bardeen1957theory ; bogoljubov1958new , slow collective modes such as critical pairing fluctuations near the metal–superconductor phase transition larkin2005theory ; schmid1966time ; maki1968critical ; cyrot1973ginzburg ; schuller2006time , or topological excitations such as vortices carrying quantized magnetic flux tinkham ; blatter1994vortices ; halperin1979resistive ; Beasley ; Nelson_VL .
-Over the years, several experimental techniques have been developed to probe these universal, low-energy features comin2016resonant ; kaindl2000ultrafast ; de2021colloquium ; parker:2010 ; essaleh:2018 ; hore:2015 .
+Yet, it remains experimentally challenging to investigate such dynamics with high spectroscopic resolution in both frequency and momentum space, particularly in two-dimensional correlated systems.
+Here, we leverage Nitrogen-Vacancy (NV) centers in diamond as a powerful and non-invasive tool to study thin-film Bi2Sr2CaCu2O8+δ (BSCCO), revealing several distinct dynamical phenomena across the superconducting phase diagram.
+At zero magnetic field and low temperatures, NV depolarization ($T_{1}$) noise spectroscopy captures the low-frequency (GHz-scale) magnetic noise generated by nodal superconducting quasiparticle excitations, in agreement with Bardeen–Cooper–Schrieffer (BCS) mean-field theory.
+Near the critical temperature $T_{c}\approx 90$ K, supercurrent-fluctuation-induced noise leads to a sharp reduction of the NV $T_{1}$.
+By carefully analyzing the temperature-scaling of $T_{1}$, we observe clear deviations from the BCS prediction, reflecting the importance of order parameter fluctuations and enabling the determination of both static and dynamical critical exponents.
+When a small field is applied, we detect a broad and asymmetric reduction of NV $T_{1}$ near $T_{c}$; the field-induced smearing of the transition unveils the presence of a vortex liquid phase.
+Finally, NV decoherence ($T_{2}$) noise spectroscopy allows us to characterize magnetic noise at even lower MHz-scale frequencies and obtain evidence for complex vortex-solid fluctuations well below $T_{c}$.
+Our results establish quantum noise spectroscopy as a versatile platform for probing dynamical phenomena in superconductors, with frequency and length scales complementary to existing techniques.
+Introduction — Strong electronic correlations are responsible for the plethora of complex emergent phenomena observed in quantum materials.
+This is perhaps best exemplified in unconventional superconductors, where electronic interactions induce a macroscopic quantum coherence of electron-pairs and dissipationless flow of current [70, 37, 75].
+Key features of this state are encoded in the low-energy dynamics, manifesting as quasiparticle excitations arising from the breaking of electron-pairs [4, 5, 10], slow collective modes such as critical pairing fluctuations near the metal–superconductor phase transition [46, 65, 53, 14, 66], or topological excitations such as vortices carrying quantized magnetic flux [70, 9, 29, 7, 58].
+Characterizing these dynamics thus offers a new avenue towards understanding the superconducting state.
+This is particularly important when addressing the emerging landscape of correlated, two-dimensional superconducting materials [3, 42], where the application of conventional experimental probes is a notorious challenge.
 Figure 1: NV noise spectroscopy of superconducting dynamics.
-(a) Summary of different experimental techniques for probing superconductivity, organized by their frequency and lengthscales, highlighting the complementarity of our qubit-based noise spectroscopy (NV $T_{1}$ and $T_{2}$) and DC magnetometry barry:2020b .
-Techniques depicted include: electrical transport essaleh:2018 , superconducting quantum interference device (SQUID) kirtley:1999 ; dusad:2019 , Lorentz microscopy harada:1992 ; silhanek:2021 , X-ray/neutron scattering comin2016resonant ; muhlbauer:2019 , magnetic resonance with nuclear, electronic, or muonic spins (NMR/ESR/µSR) hore:2015 ; hillier:2022 , angle-resolved photoemission spectroscopy (APRES) liu2008development ; kitamura2022development , low-energy and transmission electron microscopy (LEEM/TEM) altman:2010 ; moradifar:2023 , and scanning tunneling microscope (STM) parker:2010 .
+(a) Summary of different experimental techniques for probing superconductivity, organized by their frequency and lengthscales, highlighting the complementarity of our qubit-based noise spectroscopy (NV $T_{1}$ and $T_{2}$) and DC magnetometry [6].
+Techniques depicted include: electrical transport [22], superconducting quantum interference device (SQUID) [40, 20], Lorentz microscopy [30, 67], X-ray/neutron scattering [12, 57], magnetic resonance with nuclear, electronic, or muonic spins (NMR/ESR/µSR) [34, 31], angle-resolved photoemission spectroscopy (APRES) [51, 41], low-energy and transmission electron microscopy (LEEM/TEM) [2, 56], and scanning tunneling microscope (STM) [60].
 Some techniques, while capable of accessing high momentum data, require macroscopic sample volumes and, thus, are only suitable for probing mm-scale samples.
-Techniques like pump-probe spectroscopy kaindl2000ultrafast ; de2021colloquium , which are known to drive the sample strongly out of equilibrium, are not included.
+Techniques like pump-probe spectroscopy [36, 16], which are known to drive the sample strongly out of equilibrium, are not included.
 (b) Schematic of the experimental set-up: a thin film ($\sim 200$ nm thickness) of BSCCO is directly exfoliated onto a diamond plate containing an ensemble of NV centers within a thin layer ($\sim 50$ nm) beneath its surface.
@@
 The NV exhibits a splitting of its $|m_{s}=\pm 1\rangle$ sublevels, which is determined by the local magnetic field $B_{z}$, centered around a zero-field splitting $D_{\mathrm{gs}}=2.87~{}$GHz.
-(c) Schematic H-T phase diagram of a typical type-II superconductor tinkham ; keimer2015quantum ; zhou2021high .
+(c) Schematic H-T phase diagram of a typical type-II superconductor [70, 37, 75].
 The green arrows indicate zero-field cooling ($H=0$) and field-cooling ($H>0$) pathways used in our experiments.
-Broadly, techniques for measuring correlated quantum dynamics may be categorized into scattering methods, local spectroscopic probes, and bulk measurements [Fig.
+Over the past decades, several experimental techniques have been developed to probe correlated quantum dynamics in superconductors [12, 36, 16, 60, 22, 34].
+Broadly, such techniques may be categorized into scattering methods, local spectroscopic probes, and bulk measurements [Fig.
 1(a)].
 Scattering probes (e.g.
-X-ray scattering comin2016resonant ) provide both energy and momentum-resolved data, but they typically require experimental apparatus beyond table-top settings, large sample volumes, and can drive the system out of equilibrium kaindl2000ultrafast ; de2021colloquium .
-Local probes such as scanning tunneling microscopy (STM) parker:2010 are applicable to low-dimensional samples and provide lattice-scale spatial resolution, yet they exhibit limited momentum resolution and typically operate at low temperatures.
-By contrast, bulk measurements (including transport essaleh:2018 and nuclear magnetic resonance hore:2015 ) can operate over a wide temperature range, but generally demand large sample volume.
+X-ray scattering [12]) provide both energy- and momentum-resolved data, but they typically require experimental apparatus beyond table-top settings, large sample volume, and can drive the system out of equilibrium [36, 16].
+Local probes such as scanning tunneling microscopy (STM) [60] are applicable to low-dimensional samples and provide lattice-scale spatial resolution, yet they exhibit limited momentum resolution and typically operate at low temperatures.
+By contrast, bulk measurements (including transport [22] and nuclear magnetic resonance [34]) can operate over a wide temperature range, but generally demand large sample volume.
 Therefore, resolving spectral features at long wave-lengths below the meV scale in two-dimensional materials and thin film samples remains a challenge, leading to an unfulfilled gap for experimental techniques that can probe equilibrium dynamics at both low-energies and a variety of length-scales.
-In this work, we fill this gap by introducing a non-invasive, table-top experimental technique — quantum noise spectroscopy — to probe superconducting dynamics with simultaneous spatial and temporal resolution.
+In this work, we fill this gap by introducing a non-invasive, table-top experimental technique — quantum noise spectroscopy — to probe a wide range of low-energy dynamical phenomena in superconductors.
 Specifically, by using an ensemble of nitrogen-vacancy (NV) centers positioned in close proximity to a thin film high-$T_{c}$ cuprate — Bi2Sr2CaCu2O8+δ (BSCCO, Bi-2212), we perform noise spectroscopy of superconducting fluctuations as a function of temperature, frequency and applied magnetic field.
@@
 1(a)].
-Our main results are threefold.
-First, at zero applied field, we study the fluctuations arising from both long-lived quasiparticle excitations at low temperature, and collective critical modes near the metal-superconductor transition temperature $T_{c}$.
+Our results are threefold.
+First, at zero applied field, we observe and characterize the fluctuations arising from both long-lived quasiparticle excitations at low temperature, and collective critical modes near the metal-superconductor transition temperature $T_{c}$.
 These fluctuations induce GHz-scale magnetic noise at the location of the nearby NVs which can be characterized via NV depolarization ($T_{1}$) spectroscopy.
-Leveraging this capability, we quantitatively measure the characteristic time-scale $\tau_{0}$ of critical slowing down of pairing fluctuations near $T_{c}$, as well as the static and dynamical critical exponents associated with the phase transition landau2013statistical ; landau1980statistical ; hohenberg1977theory .
-Second, we directly probe the broadening of the transition in the presence of a small applied field $H$.
-At non-zero $H$, the system hosts superconducting vortices whose dynamics induce additional magnetic field fluctuations.
-The field dependence of the NV relaxation rate is consistent with the existence of an intermediate-temperature vortex liquid phase blatter1994vortices ; halperin1979resistive , where magnetic flux lines diffuse through the superconducting film.
-Finally, we employ decoherence ($T_{2}$) noise spectroscopy to study lower-frequency fluctuations in the MHz regime machado2023quantum .
+These capabilities enable us to directly measure the characteristic timescale $\tau$ of critical slowing down of pairing fluctuations near $T_{c}$ and determine both the static and dynamical critical exponents associated with the phase transition [45, 44, 33].
+Second, we directly measure the broadening of the transition in the presence of a small applied field $H$.
+At non-zero $H$, the system also hosts superconducting vortices whose dynamics induce additional magnetic field fluctuations.
+The field dependence of the NV relaxation rate is consistent with the existence of an intermediate-temperature vortex liquid phase [9, 29], where magnetic flux lines diffuse through the superconducting film.
+Finally, we employ decoherence ($T_{2}$) noise spectroscopy to study lower-frequency fluctuations in the MHz regime [52].
 Deep within the superconducting phase, we observe a distinct noise mechanism that significantly enhances the NV decoherence rate.
-These strong, low-frequency magnetic fluctuations potentially arise from thermally assisted flux motion in a vortex solid phase blatter1994vortices ; MFisher89 ; fisherfisherhuse .
+These strong, low-frequency magnetic fluctuations potentially arise from thermally assisted flux motion in a vortex solid phase [9, 24, 23].
 Figure 2: Observing Meissner effect and superconducting vortices.
@@
 The gray dashed line marks the superconducting transition temperature $T_{c}\approx 90$ K.
-Top: at 10 K, ODMR spectrum of NV centers under BSCCO at $H=0$ (yellow), where the intrinsic splitting $2\Pi_{\perp}$ originates from the local electric field mittiga2018imaging .
+Top: at 10 K, ODMR spectrum of NV centers under BSCCO at $H=0$ (yellow), where the intrinsic splitting $2\Pi_{\perp}$ originates from the local electric field [55].
 At $H=15$ G, the spectrum (blue) remains largely unchanged, indicating a strong magnetic field repulsion due to the Meissner effect.
@@
 (e) Schematic of magnetic flux being repelled by a superconductor in its Meissner phase.
-Experimental setup— For our experiments, we choose Bi2Sr2CaCu2O8+δ (Bi-2212), a prominent type-II superconductor with a layered crystal structure and a high critical temperature $T_{c}$ exceeding the boiling point of liquid nitrogen keimer2015quantum ; zhou2021high ; blatter1994vortices .
-This compound exhibits a rich and complex H-T phase diagram, making it an ideal platform for investigating a variety of different types of low-energy dynamics [Fig.
+Experimental setup — For our experiments, we study Bi2Sr2CaCu2O8+δ (BSCCO), a prominent type-II superconductor with a layered crystal structure and a high critical temperature $T_{c}$ exceeding the boiling point of liquid nitrogen [37, 75, 9].
+This compound exhibits a rich and complex phase diagram, making it an ideal platform for investigating a variety of different types of low-energy dynamics [Fig.
 1(c)].
-Bulk crystals of nearly optimally doped BSCCO-2212 are grown by the traveling floating zone method wen2008large , with $T_{c}\approx 90~{}$K measured via transport (see Methods).
+Bulk crystals of nearly optimally doped BSCCO are grown by the traveling floating zone method [71], with $T_{c}\approx 90~{}$K measured via transport (see Methods).
 To perform noise spectroscopy with NV centers, we exfoliate the BSCCO to a 200200200~{}200nm thick film and transfer it directly onto a diamond sample containing an ensemble of NV centers in a thin layer $\sim 50$ nm beneath the diamond surface [Fig.
 1(b)].
+Each NV center hosts a spin-1 electronic ground state, with Hamiltonian [63]:
+$\begin{split}\mathcal{H}=&~{}D_{\mathrm{gs}}S_{z}^{2}+\gamma_{e}B_{z}S_{z}\\ &+\Pi_{x}(S_{y}^{2}-S_{x}^{2})+\Pi_{y}(S_{x}S_{y}+S_{y}S_{x}),\end{split}$ (1)
+where {$S_{x}$, $S_{y}$, $S_{z}$} are the spin-1 operators of the NV with the N–V axis defining the quantization axis ($\hat{z}$), $D_{\mathrm{gs}}=2.87~{}$GHz is the zero-field splitting between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sublevels, $\Pi_{\perp}=\sqrt{\Pi_{x}^{2}+\Pi_{y}^{2}}=6.6~{}$MHz characterizes the coupling to local electric fields [55], $B_{z}$ is the local magnetic field along the NV axis, and $\gamma_{e}$ is the electronic spin gyromagnetic ratio.
+The NV energy levels can be probed via optically detected magnetic resonance (ODMR) spectroscopy where one detects a decrease in NV fluorescence intensity when sweeping the frequency of an applied microwave through the resonances $|m_{s}=0\rangle\leftrightarrow|m_{s}=\pm 1\rangle$.
+In the experiment, we use a [111]-cut diamond crystal and apply a tunable magnetic field $H$ along the out-of-plane direction [Fig.
+1(b)].
+In this geometry, the NV crystallographic group aligned with the applied field exhibits the largest splitting, $\Delta=\sqrt{(2\Pi_{\perp})^{2}+(2\gamma_{e}B_{z})^{2}}$ between $|m_{s}=\pm 1\rangle$ sublevels, where $B_{z}$ includes both contributions from the applied field $H$ and the $B^{\mathrm{s}}_{z}$ generated from the superconducting sample ($B_{z}=H+B_{z}^{\mathrm{s}}$).
+The other three NV groups are always observed to be degenerate.
+Imaging the Meissner effect and flux vortices — Using the NV centers as sensitive local magnetometers, we begin by determining the presence of superconductivity in BSCCO via the DC Meissner effect [64, 74, 48, 8, 59, 32].
+After zero-field cooling the sample down to 101010~{}10K (well below $T_{c}$), we perform an ODMR measurement with $H=0$ [Fig.
+2(a)].
+The resulting spectrum shows two resonances separated by the intrinsic splitting $2\Pi_{\perp}$.
+When we apply a small field $H=15~{}$G, the ODMR spectrum remains largely unchanged, displaying only a slightly larger splitting which suggests that the magnetic field generated by the sample cancels the applied one, $B_{z}^{\mathrm{s}}\approx-H$.
+Indeed, the extracted local field at the NV, $B_{z}=2.5~{}$G $\ll H=15~{}$G, corroborating the presence of superconducting Meissner phase that expels the external field.
+We emphasize that the observation of a near-complete cancellation of $H$ implies a close proximity of our NV sensors to the superconducting sample.
+As we increase the temperature, the near-complete expulsion of the external field persists until $T\sim 40~{}$K.
+Beyond this point, ODMR spectrum splits into a characteristic four resonances, indicating that the field is no longer fully canceled and that the system has transitioned into a vortex phase, allowing magnetic flux to penetrate the sample.
+As the temperature further rises, the vortex density increases, raising the local field, $B_{z}$ until it reaches the applied field $H$ at $T=T_{c}\approx 90~{}$K.
+Interestingly, right below $T_{c}$, the local field $B_{z}$ is measured to be slightly greater than $H$, which may be attributed to a paramagnetic Meissner effect [25].
+Leveraging the local nature of our sensor, we then investigate the spatial dependence of the Meissner suppression.
+Performing ODMR spectroscopy along a line-cut across the BSCCO sample [black dashed line in Fig.
+2(b)], we directly extract the local magnetic field $B_{z}$ as a function of spatial position at $T=30$ K [Fig.
+2(c)].
+Beneath the sample, the Meissner suppression is clearly observed at small applied field $H\lesssim 20~{}$G.
+However, when $H$ exceeds 25 G, which is on the same order as the independently characterized lower critical field $H_{\mathrm{c1}}$ (see Methods), the Meissner effect is suppressed, and $H$ is only partially expelled.
+Crucially, the residual field also exhibits spatial variations across the sample, indicating that magnetic flux penetrates the BSCCO, forming superconducting vortices.
+One such vortex, possibly pinned by a local defect, is imaged in Fig.
+2(d), where the locally measured magnetic field $B_{z}$ exceeds the applied field $H=40~{}$G due to the bunching of magnetic field lines.
 Figure 3: Zero-field superconducting fluctuations.
@@
 (b) The $T_{1}$ relaxation timescales as a function of temperature measured on the NV centers located under and away from BSCCO.
-(c) Spin relaxation rate, $1/T_{1}$, along a spatial line-cut across the BSCCO sample [see (a) inset] at 110, 93 and 40 K.
+(c) Spin relaxation rate, $1/T_{1}$, along a spatial line-cut across the BSCCO sample (see inset) at 110, 93 and 40 K.
 The relaxation rate increases for NV centers under BSCCO (shadow region), particularly near the critical temperature $T_{c}$.
@@
 This data aligns closely with a theoretical model of critical Langevin dynamics of the order parameter (solid lines).
-Each NV center hosts a spin-1 electronic ground state, with Hamiltonian rovny2024nanoscale :
-$\begin{split}\mathcal{H}=&~{}D_{\mathrm{gs}}S_{z}^{2}+\gamma_{e}B_{z}S_{z}\\ &+\Pi_{x}(S_{y}^{2}-S_{x}^{2})+\Pi_{y}(S_{x}S_{y}+S_{y}S_{x}),\end{split}$ (1)
-where {$S_{x}$, $S_{y}$, $S_{z}$} are the spin-1 operators with the N–V axis defining the spin-quantization axis ($\hat{z}$), $D_{\mathrm{gs}}=2.87~{}$GHz is the zero-field splitting between $|m_{s}=0\rangle$ and $|m_{s}=\pm 1\rangle$ sublevels, $\Pi_{\perp}=\sqrt{\Pi_{x}^{2}+\Pi_{y}^{2}}=6.6~{}$MHz characterizes the coupling to local electric fields mittiga2018imaging , $B_{z}$ is the local magnetic field along the NV axis, and $\gamma_{e}$ is the electronic spin gyromagnetic ratio.
-The NV energy levels can be probed via optically detected magnetic resonance (ODMR) spectroscopy where one detects a decrease in NV fluorescence intensity when sweeping the frequency of an applied microwave through the resonances $|m_{s}=0\rangle\leftrightarrow|m_{s}=\pm 1\rangle$.
-In the experiment, we use a [111]-cut diamond crystal and apply a tunable magnetic field $H$ applied along the out-of-plane direction [Fig.
+Zero-field superconducting fluctuations — Employing the full capabilities of the NV center, we go beyond conventional static magnetic field measurements and study superconducting fluctuations by performing $T_{1}$ relaxometry of the NV [Fig.
 1(b)].
-In this geometry, the NV crystallographic group aligned with the applied field exhibits the largest splitting, $\Delta=\sqrt{(2\Pi_{\perp})^{2}+(2\gamma_{e}B_{z})^{2}}$ between $|m_{s}=\pm 1\rangle$ sublevels, where $B_{z}$ includes both contributions from the applied field $H$ and the $B^{\mathrm{s}}_{z}$ generated from the superconducting sample ($B_{z}=H+B_{z}^{\mathrm{s}}$).
-The other three NV groups are always observed to be degenerate.
-Imaging the Meissner effect and flux vortices — Using the NV centers as sensitive local magnetometers, we begin by determining the presence of superconductivity in BSCCO via the DC Meissner effect schlussel2018wide ; yip2019measuring ; lesik2019magnetic ; bhattacharyya2023imaging ; nusran2018spatially ; ho2024studying .
-After zero-field cooling the sample down to 101010~{}10K (well below $T_{c}$), we perform an ODMR measurement with $H=0$ [Fig.
-2(a)].
-The resulting spectrum shows two resonances separated by the intrinsic splitting $2\Pi_{\perp}$.
-When we then apply a small field $H=15~{}$G, the ODMR spectrum remains largely unchanged, displays only a slightly larger splitting, suggesting that the magnetic field generated by the sample, $B_{z}^{\mathrm{s}}\approx-H$.
-Indeed, the extracted local field at the NV, $B_{z}=2.5~{}$G $\ll H=15~{}$G, corroborating the presence of superconducting Meissner phase that expels the external field.
-We emphasize that the observation of a near-complete cancellation of $H$ underscores the close proximity of our NV sensors to the superconducting BSCCO.
-As we increase the temperature, the near-complete expulsion of the external field persists $T\sim 40~{}$K.
-Beyond this point, ODMR spectrum splits into a characteristic four resonances, indicating that the field is no longer fully canceled and that the system has transitioned into a vortex phase, allowing magnetic flux to penetrate the sample.
-As the temperature further rises, the vortex density increases, raising the local field, $B_{z}$ until it reaches the applied field $H$ at $T=T_{c}\approx 90~{}$K, Interestingly, right below $T_{c}$, the local field $B_{z}$ is measured to be slightly greater than $H$, which may be attributed to a paramagnetic Meissner effect geim1998paramagnetic .
-Leveraging the local nature of our sensor, we then investigate the spatial dependence of the Meissner suppression.
-Performing ODMR spectroscopy along a line-cut across the BSCCO [black dashed line in Fig.
-2(b)], we directly extract the local magnetic field $B_{z}$ as a function of spatial position at $T=30$ K [Fig.
-2(c)].
-Beneath the sample, the Meissner suppression is clearly observed at small applied field $H\lesssim 20~{}$G.
-However, when $H$ exceeds 25 G, which is on the same order as the independently characterized lower critical field $H_{\mathrm{c1}}$ (see Methods), the Meissner effect is suppressed, and $H$ is only partially expelled.
-Crucially, the residual field also exhibits spatial variations across the sample, indicating that magnetic flux penetrates the BSCCO, forming superconducting vortices.
-One such vortex, possibly pinned by a local defect, is imaged in Figure 2(d), where the locally measured magnetic field $B_{z}$ exceeds the applied field $H=40~{}$G due to the bunching of magnetic field lines.
-Zero-field superconducting fluctuations — Employing the full capabilities of the NV center, we go beyond conventional static magnetic field measurements by performing $T_{1}$ relaxometry of the NV to study superconducting fluctuations [Fig.
-1(a)].
 To this end, we first optically polarize the NV spin state to $|m_{s}=0\rangle$ and subsequently measure the timescale $T_{1}$ over which the spin relaxes back to thermal equilibrium [Fig.
 3(a)].
-Although the intrinsic relaxation of NV centers is primarily driven by spin-phonon interactions within the diamond itself, additional magnetic fluctuations from BSCCO, resonant with the NV’s bare frequency ($D_{\mathrm{gs}}=2.87~{}$GHz), can expedite the spin relaxation process and reduce the $T_{1}$ of NV.
+Although spin-phonon interactions within the diamond itself lead to an intrinsic relaxation rate of NV centers, additional magnetic fluctuations from BSCCO, resonant with the NV’s bare frequency ($D_{\mathrm{gs}}=2.87~{}$GHz), can expedite the spin relaxation process and reduce the $T_{1}$ of NV.
 This is indeed borne out by our data.
-As shown in Figures 3(a)(b), NV centers located beneath the BSCCO exhibit shorter $T_{1}$ compared to NV centers situated away from the BSCCO across the entire temperature range covering the superconducting phase, criticality and metallic phase.
-This trend is further highlighted in Figure 3(c), which shows that the NV relaxation rate ($1/T_{1}$) as a function of a spatial line-cut across three distinct temperature regimes, with faster relaxation rates observed in all cases for NV centers beneath the BSCCO sample.
+As shown in Figures 3(a)(b), NV centers located beneath the BSCCO exhibit shorter $T_{1}$ compared to NV centers situated away from the BSCCO across the entire temperature range.
+This trend is further highlighted in Figure 3(c), which shows that the NV relaxation rate ($1/T_{1}$) as a function of a spatial line-cut across three distinct temperature regimes — the superconducting phase, the critical region and the metallic phase, with faster relaxation rates observed in all cases for NV centers beneath the BSCCO sample.
 To quantitatively analyze the observed magnetic noise, we extract the BSCCO-induced decay rate, $\Gamma_{1}(T)$, by subtracting the intrinsic relaxation rate — measured from NV centers located far from the BSCCO sample — from the total relaxation rate of NV centers situated beneath the BSCCO [Fig.
 3(d)].
-At high temperatures, $T>T_{c}$, $\Gamma_{1}(T)$ is finite due to Johnson-Nyquist noise in the metallic phase kolkowitz2015probing ; agarwal2017magnetic ; hsieh2019imaging .
-At very low temperatures, $T\ll T_{c}$, the magnetic noise is significantly suppressed in the superconductor.
-Crucially, the form of this suppression offers insights into the nature of the pairing symmetry.
-While in conventional s-wave superconductors $\Gamma_{1}(T)$ is exponentially suppressed at low temperatures due to a fully gapped Bogoliubov quasiparticle spectrum  bardeen1957theory ; bardeen1957microscopic ; chatterjee2022single ; dolgirev2022characterizing ; kelly2024superconductivity ; li2024observation , the presence of nodal quasiparticles in d-wave layered superconductors implies a distinct power-law suppression of $\Gamma_{1}(T)\propto T^{2}$ dolgirev2022characterizing ; chatterjee2022single .
+At high temperatures, $T>T_{c}$, Johnson-Nyquist noise in the metallic phase determines the strength of the magnetic noise [43, 1, 35].
+At very low temperatures, $T\ll T_{c}$, the magnetic noise is significantly suppressed.
+Crucially, the form of this suppression offers insights into the nature of the superconducting pairing symmetry.
+While in conventional s-wave superconductors $\Gamma_{1}(T)$ is exponentially suppressed at low temperatures due to a fully gapped Bogoliubov quasiparticle spectrum  [5, 4, 11, 18, 38, 49], the presence of nodal quasiparticles in d-wave layered superconductors implies a distinct power-law suppression of $\Gamma_{1}(T)\sim T^{2}$ [18, 11].
 Indeed, this expectation is consistent with our observations [Fig.
-3(d), inset], corroborating the nodal d-wave nature of superconductivity in BSCCO keimer2015quantum .
+3(d), inset], corroborating the nodal d-wave nature of superconductivity in BSCCO [37].
 The most striking feature of Fig.
-3(d) is the sharp and symmetric peak in $\Gamma_{1}$ near the critical temperatures $T_{c}$.
+3(d) is the sharp and symmetric peak in $\Gamma_{1}(T)$ near the critical temperatures $T_{c}$.
 This peak arises from strong fluctuations in the superconducting order parameter, resulting in enhanced current fluctuations that act as source of magnetic noise at the metal-superconductor phase transition.
-Crucially, this enhancement is observed across the entire BSCCO sample, as emphasized by the spatial line cut near $T_{c}$, Fig.
-3(c).
+Crucially, this enhancement is observed across the entire BSCCO sample [Fig.
+3(c)].
 Interestingly, near $T_{c}$, $\Gamma_{1}(T)$ begins to show marked deviations from mean-field expectations for d-wave superconductors.
-In particular, mean-field theory predicts a divergence of $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$ up to logarithmic corrections dolgirev2022characterizing .
-In contrast, the experimentally measured $\Gamma_{1}$ at criticality is distinctly more singular, with $x\approx 1$ [Fig.
+In particular, mean-field theory predicts a divergence of $\Gamma_{1}(T)\propto|T-T_{c}|^{-x}$ with $x=1/2$ (up to logarithmic corrections) [18].
+By contrast, the experimentally measured $\Gamma_{1}(T)$ at criticality is distinctly more singular, with $x\approx 1$ [Fig.
 3(e)].
-This deviation arises because BCS theory neglects both amplitude and phase fluctuations of the superconducting order parameter $\psi$, which play a significant role near the critical point emery1995importance ; maki1968critical .
-These effects are particularly important for thin-film superconductors where fluctuations are expected to be stronger tinkham .
-Consequently, quantitative characterizing the divergent time-scale associated with dissipative dynamics of the superconducting order parameter near $T_{c}$ remains an important open question, which we turn to next.
-To properly account for low-energy critical fluctuations, we consider the following phenomenological Langevin dynamics of the superconducting order parameter $\psi(\mathbf{r},t)$.
+This deviation arises because BCS theory neglects both amplitude and phase fluctuations of the superconducting order parameter $\psi$, which play a significant role near the critical point [21, 46, 53].
+These effects are particularly important for thin-film superconductors where fluctuations are expected to be stronger [70].
+Consequently, a quantitative characterization of the divergent time-scale associated with dissipative dynamics of the superconducting order parameter near $T_{c}$ remains an important open question, which we turn to next.
+To properly account for low-energy critical fluctuations, we consider the following phenomenological Langevin dynamics of the superconducting order parameter $\psi(\mathbf{r},t)$:
 $\partial_{t}\psi(\mathbf{r},t)$ $=-\gamma\frac{\delta F}{\delta\Psi^{*}(\mathbf{r},t)}+\eta(\mathbf{r},t),$
 […]
-where $F[\psi]$ is the Ginzburg-Landau free energy functional landau1980statistical ; cyrot1973ginzburg ; schuller2006time ; tinkham , $\eta$ is local white noise that captures the effect of coarse-graining, and $\gamma$ sets the rate of relaxation towards equilibrium hohenberg1977theory .
-Crucially, by setting $r(T)=a(T-T_{c})$ with $a>0$, we can analyze fluctuations of the order parameter about the minima of the free energy: $\langle\psi\rangle=0$ for $T>T_{c}$ (metallic) and $\langle\psi\rangle=\psi_{0}$ for $T<T_{c}$ (superconducting).
-Such fluctuations in turn leads to enhanced current-induced magnetic noise at the NV frequency $D_{\mathrm{gs}}$, enhancing the relaxation rate $\Gamma_{1}(T)$ near $T_{c}$.
-This enhancement of the noise spectral weight at low frequency is a consequence of critical slowing down of order parameter fluctuations: for weak-coupling superconductors it is quantified by the diverging timescale $\tau_{\rm GL}=8\hbar/(k_{B}|T-T_{c}|)$ larkin2005theory ; schmid1966time near $T_{c}$.
-We find that our theoretical model (see Methods) can faithfully reproduce both the magnitude of the magnetic noise, and the divergence of $\Gamma_{1}(T)$ near $T_{c}$ with the experimentally observed power-law $x=1$ on both sides of the transition [Fig.
+where $F[\psi]$ is the Ginzburg-Landau free energy functional [44, 14, 66, 70], $\eta$ is local white noise that captures the effect of coarse-graining, and $\gamma$ sets the rate of relaxation towards equilibrium [33].
+Crucially, by setting $r(T)\propto T-T_{c}$ near the critical point, we can analyze fluctuations of the order parameter around the minima of free energy associated with both phases: $\langle\psi\rangle=0$ for $T>T_{c}$ (metallic) and $\langle\psi\rangle=\psi_{0}$ for $T<T_{c}$ (superconducting).
+These order parameter fluctuations manifest themselves as current fluctuations in the BSCCO sample which induce magnetic noise at the NV’s frequency $D_{\mathrm{gs}}$, enhancing its relaxation rate $\Gamma_{1}(T)$.
+This low-frequency fluctuation enhancement is directly connected to the phenomena of critical slowing down near the phase transition; the noise spectral density at low frequencies scales with the scattering timescale which diverges as one approaches the critical point.
+For weak-coupling superconductors, this divergent timescale is quantified by $\tau_{\rm GL}=8\hbar/(k_{B}|T-T_{c}|)$ near $T_{c}$ [46, 65].
+We find that our theoretical model (see Methods) can faithfully reproduce the divergence of $\Gamma_{1}(T)$ on both sides of the transition [Fig.
 3(e)].
-In addition, by fitting $\Gamma_{1}(T)$ with our model, we extract a quantitative time-scale for critical slowing down $\tau_{0}^{\rm fitted,M}\approx 1.2\,\tau_{\rm GL}$ on the metallic side, and $\tau_{0}^{\rm fitted,SC}\approx 0.2\,\tau_{\rm GL}$ on the superconducting side.
-Remarkably, our results indicate that the decay rates of the order parameter near criticality are comparable to the analytical weak-coupling results $\tau_{0}^{\rm wc,M}=\tau_{\rm GL}$ larkin2005theory on the metallic side and $\tau_{0}^{\rm wc,SC}=0.5\,\tau_{\rm GL}$ on the superconducting side schmid1966time ; KimDiessel , even though high $T_{c}$ cuprates like BSCCO are expected to be strong-coupling superconductors lee2006doping .
-Further, our modeling of critical noise also allows us to extract the static critical exponent $\nu$ that determines the divergence of the order parameter correlation length $\xi(T)\propto|T-T_{c}|^{-\nu}$, and the dynamical critical exponent $z$ that relates the correlation time $\tau_{0}$ and correlation length as $\tau_{0}(T)\propto\xi^{z}$.
-The close agreement of the experimental observations with Langevin dynamics of $\psi$ implies that $\nu=1/2$, and $\tau_{0}(T)\propto|T-T_{c}|^{-1}\propto\xi^{2}$ implies $z=2$, consistent with expectations for a non-conserved order parameter hohenberg1977theory .
+At the same time, our approach enables us to quantitatively extract the value of the scattering timescale $\tau^{\text{fit}}$ by fitting our model to the experimental data on both the metallic [M] and superconducting [SC] sides of the transition.
+Remarkably, the extracted value for the scattering timescale, given by $\tau^{\rm fit}_{\rm M[SC]}/\tau_{\rm GL}\approx 1.2[0.2]$, are comparable to the analytical weak-coupling results, $\tau^{\rm wc}_{\rm M[SC]}/\tau_{\rm GL}=1[0.5]$ [46, 65, 39].
+Understanding the broad agreement between the weak-coupling prediction and the expected strong-coupling nature of BSCCO superconductivity remains an open question [47].
+The agreement between our theoretical model and the experimental data suggests that the critical behavior is properly captured by our Langevin dynamics approach.
+Indeed, the resulting value of the associated critical exponents, $\nu=1/2$ and $z=2$ [33], directly reproduce the aforementioned divergence of the scattering timescale $\tau$ via the scaling behavior of the correlation length $\xi$: $\tau\propto\xi^{z}\propto|T-T_{c}|^{-z\nu}$.
 Figure 4: In-field criticality and vortex dynamics.
 (a) BSCCO-induced relaxation rate, $\Gamma_{1}(T)$ (normalized), as a function of temperature at $H=0,40,200$ G.
-Inset: the measured peak magnitude of $\Gamma_{1}$ as a function of applied field, which is consistent with a linear dependence on the field strength $H$ (dashed line).
+Inset: the measured peak magnitude of $\Gamma_{1}(T)$ as a function of applied field, which is consistent with a linear dependence on the field strength $H$ (dashed line).
 (b) BSCCO-induced decoherence rate, $\Gamma_{2}(T)$, as a function of temperature at $H=55,120,200$ G.
 Inset: The measured $T^{\mathrm{XY8}}_{2}$ decoherence timescale as a function of temperature for NVs under and away from BSCCO at $H=200$ G.
-(c) Schematic of the noise spectrum arising from different dynamical processes in BSCCO.
-The shaded area marks the most sensitive frequencies of NV relaxation ($\Gamma_{1}$) spectroscopy and decoherence ($\Gamma_{2}$) spectroscopy.
-Since the NV’s intrinsic decoherence rate far exceeds its relaxation rate, $\Gamma_{2}$ is less sensitive to weak magnetic noise than $\Gamma_{1}$.
+(c) Schematic of the noise spectrum arising from different dynamical processes in BSCCO in the context of our two probing techniques, NV relaxation ($T_{1}$) and decoherence ($T_{2}$) noise spectroscopy.
+While less sensitive than $T_{1}$ spectroscopy, decoherence spectroscopy is able to access fluctuations at much lower frequencies and observe vortex solid dynamics.
 (d) Left: a trapped vortex imaged in BSCCO under $H=40$ G at 50 K.
 Right: the measured relaxation time $T_{1}$ and decoherence time $T^{\mathrm{XY8}}_{2}$ of NV centers at different spatial positions across the vortex.
-In-field criticality and vortex dynamics — Next we explore how an out-of-plane magnetic field impacts critical fluctuations.
-Specifically, we apply a small bias magnetic field and perform $T_{1}$ noise spectroscopy at two distinct fields, $H=40$ G and $H=200$ G.
-Intriguingly, compared with the zero-field $\Gamma_{1}$, two distinct new features emerge [Fig.
-4(a)]: (i) The peak in the measured noise, $\Gamma_{1}$, is shifted towards a slightly lower temperature.
-(ii) While the noise signal approaching criticality from the metallic side remains relatively sharp, from the superconducting side, it increases much more gradually and smoothly towards criticality, resulting in a broad and asymmetric spectrum.
-The former is expected, as a perpendicular field hinders phase coherence and lowers $T_{c}$.
-The latter can be attributed to a significantly larger fluctuation regime, dominated by the two-dimensional XY critical behavior Beasley ; halperin1979resistive ; curtis2024probing .
-Below $T_{c}$, an intermediate vortex liquid phase blatter1994vortices ; fisherfisherhuse ; Nelson_VL is expected to form.
-In this phase, the motion of the superconducting vortices, which trap magnetic flux quanta, can generate strongly fluctuating magnetic fields at the NV, resulting in enhanced noise.
-Further evidence for this scenario can be gleaned from the scaling of the peak magnitude of $\Gamma_{1}$ with the magnetic field $H$ [inset, Fig.
+In-field criticality and vortex dynamics — Next we turn on the applied external field and explore the in-field critical fluctuations at $H=40~{}$G and $H=200~{}$G.
+Intriguingly, the measured $\Gamma_{1}(T)$ differs from the zero-field case in two key ways [Fig.
+4(a)]: (i) The peak position of the noise is slightly shifted towards lower temperatures.
+(ii) The peak becomes broad and highly asymmetric; while the noise signal remains sharp when approaching criticality from the metallic side, it increases much more gradually and smoothly from the superconducting side.
+The former is expected, as a perpendicular field hinders phase coherence and lowers the critical temperature $T_{c}$.
+The latter can be attributed to two different factors.
+On the one hand, an external magnetic field broadens the temperature range where critical fluctuations are important as described by two-dimensional XY critical dynamics [7, 29, 13].
+On the other hand, below $T_{c}$, an intermediate vortex liquid phase is expected to form where the local motion of the superconducting vortices (which trap magnetic flux quanta) can generate an additional fluctuating magnetic field at the NV [9, 23, 58].
+Further evidence for the latter can be gleaned from the scaling of the peak magnitude of $\Gamma_{1}$ with the magnetic field $H$ [inset, Fig.
 4(a)], where the density of vortices $n_{v}$ increases with increasing field strength.
 Within a simple model of diffusive vortex motion (see Methods), the local magnetic noise, measured via $\Gamma_{1}(H)$, scales as $D_{v}^{-1}$ where $D_{v}\sim v^{2}\tau_{v}$ is the vortex diffusion constant.
-We expect the vortex scattering time $\tau_{v}$ to scale inversely proportional to the vortex density $n_{v}$, and the mean-square speed $\langle v^{2}\rangle$ is primarily determined by thermal energy.
-Consequently, the measured magnetic noise scales as $\Gamma_{1}(H)\propto n_{v}\propto H$, in line with our observation [inset, Fig.
+The vortex scattering time $\tau_{v}$ is expected to be inversely proportional to both the vortex density $n_{v}$, and the mean-square speed $\langle v^{2}\rangle\sim T$ is primarily determined by the temperature of the sample.
+Consequently, the measured magnetic noise scales as $\Gamma_{1}(H)\propto n_{v}\propto H$, in agreement with our observations [inset, Fig.
 4(a)].
-As the temperature decreases further, the vortices either freeze into a lattice or get pinned by defects, resulting in a decrease of $\Gamma_{1}(T)$.
+As the temperature further decreases, the vortices either freeze into a lattice or become pinned by defects, resulting in a decrease of $\Gamma_{1}(T)$.
 This conclusion is consistent with our measurement of noise over a line-cut on a large pinned vortex at low temperatures shown in Fig.
-4(d) — $T_{1}$ does not change appreciably across the vortex, indicating that pinned vortices do not contribute to magnetic noise in the GHz frequency range.
-$T_{2}$ noise spectroscopy of slow vortex dynamics — At low temperatures, one enters deep into the superconducting phase and the vortex motion freezes out.
-Nevertheless, with $H>H_{c1}$, BSCCO can exhibit complex vortex dynamics, such as thermally assisted flux jumps between different pinning centers, or de-pinning of vortices to exit the superconducting sample, potentially leading to a revival of magnetic noise.
-However, our data for $\Gamma_{1}(T)$ in the low temperature regime exhibits no additional features at finite $H$.
-One possibility is that, unlike the metal-superconducting transition, fluctuations arising from vortex dynamics at low temperatures occur at a much lower frequency range, and therefore eludes detection by $T_{1}$ spectroscopy that is only sensitive to the noise spectrum in the GHz range.
-To this end, we perform NV decoherence ($T_{2}$) noise spectroscopy to capture the magnetic noise spectrum in the MHz frequency range machado2023quantum ; xue2024signatures ; ziffer2024quantum ; li2024critical .
-Specifically, after preparing the NV center into a quantum superposition state, $(|m_{s}=0\rangle+|m_{s}=-1\rangle)/\sqrt{2}$, we apply an advanced dynamical decoupling sequence (XY-8, see Supplementary Information) to measure the decoherence rate induced by BSCCO, $\Gamma_{2}(T)$, as a function of both temperature and bias external field $H$ [Fig.
+4(d) — $T_{1}$ does not change appreciably across the vortex, indicating that pinned vortices do not exhibit dynamics in the GHz frequency range.
+$T_{2}$ noise spectroscopy of slow vortex dynamics — At sufficiently low temperatures, the system is deep in the superconducting phase and the vortex liquid is expected to transition into a solid state.
+Nevertheless, with $H>H_{c1}$, BSCCO can still exhibit complex vortex dynamics, such as thermally assisted flux jumps between different pinning centers, or the de-pinning of vortices to exit the superconducting sample, leading to fluctuating magnetic signals from different physical processes.
+However, the measured $\Gamma_{1}(T)$ at low temperature regime does not display any features, suggesting that these processes operate at even lower frequency, and therefore elude detection by $T_{1}$ spectroscopy sensitive only to GHz-range fluctuations [Fig.
+4(c)].
+To investigate these dynamics, we perform NV decoherence ($T_{2}$) noise spectroscopy to capture the magnetic noise spectrum in the MHz frequency range [52, 73, 76, 50].
+Specifically, after preparing the NV center into a quantum superposition state, $(|m_{s}=0\rangle+|m_{s}=-1\rangle)/\sqrt{2}$, we apply a dynamical decoupling sequence (XY-8, see Supplementary Information) to measure the decoherence rate induced by BSCCO, $\Gamma_{2}(T)$, as a function of both temperature and applied field $H$ [Fig.
 4(b)].
 Remarkably, $\Gamma_{2}(T)$ exhibits a broad noise spectrum with several prominent features.
-(i) The onset temperature for $\Gamma_{2}(T)$ progressively decreases with increasing field $H$, and does not coincide with the peak of $\Gamma_{1}$ near criticality.
-(ii) For a fixed $H$, $\Gamma_{2}(T)$ reaches its maximum value at a much lower temperature $T^{*}$ than $\Gamma_{1}(T)$; however, $H$ is larger than the independently measured $H_{c1}$ of the sample at $T^{*}$ (see Methods), indicating that vortices are still present in the sample.
-(iii) The maximum value of $\Gamma_{2}(T)$ is nearly independent of the external field $H$.
-To understand this behavior, we begin by noting that the decoherence time $T_{2}$ on and off BSCCO is nearly identical in both the metallic phase and near criticality [inset of Fig.
+(i) With increasing field $H$, there is a reduction of the onset temperature for $\Gamma_{2}$-type noise and this onset occurs for temperatures below the peak position of $\Gamma_{1}(T)$.
+(ii) For different $H$, $\Gamma_{2}(T)$ reaches its maximum value at a much lower temperature $T^{*}$ than $\Gamma_{1}(T)$; however, the applied field $H$ remains larger than the independently measured $H_{c1}$ of the sample at $T^{*}$ (see Methods), indicating that vortices are still present in the sample.
+(iii) The maximum value of $\Gamma_{2}(T)$ is insensitive to the strength of the external field $H$.
+To understand these observations, we must rule some of the previously studied mechanisms.
+In particular, we begin by noting that the decoherence time $T_{2}$ on and off BSCCO is nearly identical in both the metallic phase and near criticality [inset of Fig.
 4(b)].
 This observation indicates that $T_{2}$ noise spectroscopy is less sensitive than $T_{1}$ in detecting both Johnson-Nyquist noise and critical current fluctuations.
-This difference in sensitivity can be attributed to the intrinsic timescale gap between $T_{1}$ and $T_{2}$.
-Specifically, the NV $T_{1}$ is around three orders of magnitude larger than $T_{2}$, offering a much longer time window to accumulate the noise signals (see Methods).
-Consequently, a significantly stronger noise amplitude from BSCCO is required to induce a measurable $T_{2}$ in NV sensors.
-Furthermore, $T_{2}$ measured across a line-cut intersecting a pinned vortex show no position dependence, indicating that frozen vortices also do not contribute to magnetic noise in the MHz frequency range.
-Combining these observations, we deduce that $T_{2}$ is most sensitive to vortex motion within the vortex solid phase MFisher89 ; fisherfisherhuse .
-The substantial noise magnitude suggests that it likely originates from thermal fluctuations, which can cause positional jumps of fluxes or flux bundles between different pinning centers, resulting in sudden large changes in the local magnetic field.
+This difference in sensitivity can be attributed to the separation between the intrinsic depolarization and dephasing timescales of the NV.
+More specifically, the NV’s $T_{1}$ is around three orders of magnitude larger than its $T_{2}$; a significantly stronger noise amplitude is required to induce a measurable change of NV $T_{2}$ (Methods).
+At the same time, $T_{2}$ measured across a line-cut intersecting a strongly pinned vortex shows no position dependence, indicating that static vortices also do not contribute to magnetic noise in the MHz frequency range.
+This suggests that $T_{2}$ is most sensitive to vortex motion within the vortex solid phase [24, 23].
+More specifically, the noise magnitude implies that it likely originates from thermal fluctuations, which can cause positional jumps of fluxes or flux bundles between different weak pinning centers, resulting in sudden large changes in the local magnetic field.
 These dynamics are relatively slow, extending only up to the MHz range.
@@
 Only at very low temperatures, approximately corresponding to the Meissner phase, does such flux motion begin to freeze out, leading to a decrease in $\Gamma_{2}(T)$ below $40$ K.
-Summary and Outlook — For the first time, we demonstrated the use of NV centers as nano-scale quantum sensors to probe three distinct aspects of low-energy dynamics in high-$T_{c}$ superconductors: gapless quasiparticle excitations at low temperatures, critical superconducting fluctuations near $T_{c}$, and vortex motion in an intermediate temperature range in presence of an external magnetic field.
-The nanoscale proximity and high sensitivity of our NV sensors, combined with their operational frequency range spanning three decades (from MHz to GHz) and sub-micron spatial resolution, enable the measurement of local dynamics beyond the capabilities of conventional techniques [Fig.
+Outlook — For the first time, we demonstrate the use of NV centers as nano-scale quantum sensors of superconducting dynamics.
+The nano-scale proximity and high sensitivity of our NV sensors, combined with their operational frequency range spanning three decades (from MHz to GHz) and sub-micron spatial resolution, enable the measurement of a wide swath of dynamical phenomena beyond the capabilities of conventional techniques [Fig.
 1(a)].
 Looking forward, our work opens the door to several intriguing directions.
-First, while our current studies involve NV center ensembles with a spread of sensor-sample distances, an immediate future direction is to precisely adjust the distance $d$ between the quantum sensor and the target material.
-This can be achieved by using NVs on a scanning tip maletinsky2012robust ; degen2008scanning ; pelliccione2016scanned , or by creating spin sensors in a given layer of Van der Waals materials such as hexagonal boron nitride gottscholl2020initialization ; gong2023coherent ; stern2022room ; gong2023isotope .
-By tuning the distance $d$, we can adjust the momentum filter function to enable controlled probing over a variety of length scales (from nanometers to micrometers)  machado2023quantum , which is crucial for imaging inhomogeneous dynamics in superconductors bhattacharyya2023imaging ; Dailledouze2025 .
-Moreover, our measurement of the critical slowing down time-scale $\tau_{0}$ places strong constraints on theoretical models of the transition and provides new opportunities to study the mechanisms for order-parameter relaxation near criticality.
-In fact, a reasonable agreement with the weak coupling result $\tau_{0}=\tau_{\rm GL}$ on the metallic side larkin2005theory , but a stronger deviation on the superconducting side schmid1966time , indicates the necessity for further microscopic investigations of pair dynamics in unconventional superconductors.
+First, while our current studies involve ensemble NV centers with a spread of sensor-sample distances, an immediate future direction is to precisely adjust the distance $d$ between the quantum sensor and the target material.
+This can be achieved by using single NV center on a scanning tip [54, 17, 61], or by creating spin sensors in a given layer of Van der Waals materials such as hexagonal boron nitride [28, 27, 68, 26].
+By tuning the distance $d$, we can adjust the momentum filter function to enable controlled probing over a variety of length scales (from nanometers to micrometers)  [52], which is crucial for imaging inhomogeneities in superconductors [8, 15].
+Moreover, our measurement of the critical slowing down time-scale $\tau$ places strong constraints on theoretical models of the transition and provides new opportunities to study the mechanisms for order-parameter relaxation near criticality.
+In fact, a reasonable agreement with the weak coupling result $\tau=\tau_{\rm GL}$ on the metallic side [46], but a stronger deviation on the superconducting side [65], indicates the necessity for further microscopic investigations of critical dynamics in unconventional superconductors.
 Finally, our technique is not limited to BSCCO, and can be readily extended to a variety of superconducting materials.
-For instance, certain superconductors such as hydrides pickard2020superconducting ; drozdov2015conventional and nickelates sun2023signatures ; wen2024probing undergo superconducting transitions under high pressure.
-Thus, moving beyond temperature-driven transitions, the integration of NV centers into high-pressure diamond anvil cells offers a pathway to study how pressure influences critical fluctuations and vortex dynamics hsieh2019imaging ; lesik2019magnetic ; yip2019measuring ; bhattacharyya2023imaging ; Dailledouze2025 .
-Acknowledgements: We gratefully acknowledge discussions with Pavel Dolgirev, Chris Laumann, Peng Cai, Panyu Hou, Srinivas Mandyam, and Zhipan Wang.
+For instance, certain superconductors such as hydrides [62, 19] and nickelates [69, 72] undergo superconducting transitions under high pressure.
+Thus, moving beyond temperature-driven transitions, the integration of NV centers into high-pressure diamond anvil cells offers a pathway to study how pressure influences critical fluctuations and vortex dynamics [35, 48, 74, 8, 15].
+Acknowledgements: We gratefully acknowledge discussions with Pavel Dolgirev, Chris Laumann, Paul Ching-Wu Chu, Liangzi Deng, Andrew Mounce, Ania Jayich, Peng Cai, Panyu Hou, Srinivas Mandyam, Zhipan Wang.
 This work is supported by NSF NRT LinQ 2152221, NSF ExpandQISE 2328837, and the Center for Quantum Leaps at Washington University.